# Fast, robust and laser-free universal entangling gates for trapped-ion quantum computing

### v1

@@ Abstract
A novel two-qubit entangling gate for RF-controlled trapped-ion quantum processors is proposed theoretically and demonstrated experimentally. The speed of this gate is an order of magnitude higher than that of previously demonstrated two-qubit entangling gates in static magnetic field gradients. At the same time, the phase-modulated field driving the gate, dynamically decouples the qubits from amplitude and frequency noise, increasing the qubits’ coherence time by two orders of magnitude. The gate requires only a single continuous RF field per qubit, making it well suited for scaling a quantum processor to large numbers of qubits. Implementing this entangling gate, we generate the Bell states |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ in $\leq 313$ $\mathrm{\mathrm{\mu}}$s with fidelities up to $98^{+2}_{-3}\leavevmode\nobreak$in a static magnetic gradient of only 19.09 T/m. At higher magnetic field gradients, the entangling gate speed can be further improved to match that of laser-based counterparts.
@@ I Introduction
Trapped atomic ions are a physical platform well suited for quantum information processing [1, 2]. Intense research into this platform has lead to outstanding performance of trapped-ion quantum processors using laser-controlled ionic qubits [3, 4, 5, 6].
Trapped ions controlled by radio frequency (RF) signals are particularly suited for scaling up quantum computers, since technological challenges associated with using laser light for coherent control of qubits are avoided in this laser-free approach [7, 8, 9, 10, 11]. With RF-controlled ions high fidelity single- and two-qubit gates have been achieved [12, 13, 14, 15, 16], as well as low cross-talk suitable for fault-tolerant quantum computing [17]. Also, complete quantum algorithms were implemented using RF-controlled ions [18, 19]. So far, however, two-qubit gate speed has lagged behind its laser-based counterparts [15]. Here, we introduce and experimentally realize a novel two-qubit entangling gate with gate speed an order of magnitude faster than previous RF-gates in static magnetic field gradients. Another feature of this gate is its built-in robustness against external noise making additional dynamical decoupling superfluous. Furthermore, only a single modulated RF-driving field per ion is required for its operation, further simplifying the scaling up of trapped-ion quantum computers.
@@ II Experimental overview
We investigate double dressed state dynamics of a single qubit and implement conditional quantum dynamics with two qubits using ${}^{171}$${}^{+}$ ions trapped in a macroscopic linear Paul trap with radial and axial trapping frequencies of $2\pi\times 380$ kHz and $\nu=2\pi\times 98.08$ kHz, respectively. The experimental setup is similar to [16] and only briefly described here. The two qubits are encoded in hyperfine states of the electronic ground state of ${}^{171}$${}^{+}$ ions, $|0\rangle\equiv{|^{2}\rm S_{1/2},\rm F=0,\rm m_{\rm F}=0\rangle}$ and $|1\rangle\equiv{|^{2}\rm S_{1/2},\rm F=1,\rm m_{\rm F}=-1\rangle}$. A static magnetic field gradient of $19.09(1)$ T/m applied along the trap axis (z-axis) yields individual Zeeman-shifts of the two qubits and results in individual resonance frequencies near $2\pi\times 12.6$ GHz, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$, of the magnetic dipole transition between qubit states. Thus, qubits are individually coherently controlled with low cross-talk using global RF radiation near $2\pi\times 12.6$ GHz [17]. The entangling gate proposed and implemented here takes advantage of a state selective force induced by the static magnetic gradient field coupling the internal qubit states to the axial vibrational states of the 2-ion crystal [7]. Therefore, when using magnetic gradient induced coupling (MAGIC), laser light is not required for implementing conditional gates with trapped ions.
The ions are cooled close to their motional ground state in two stages. Initial Doppler cooling is followed by RF sideband cooling of both present vibrational modes giving a mean phonon number of 0.6(5) in the center-of-mass (COM) mode measured with sideband thermometry [20]. The heating rate of one trapped ion in the current experimental setup is 0.19(3) phonons/ms for this mode.
We use an Arbitrary Waveform Generator (AWG) to generate the phase-modulated RF driving field that double dresses both qubits, as in[21, 22, 23]. The driving field’s action on the qubits is described by the Hamiltonian
$H_{D}=\sum_{j=1,2}$ $\Omega_{1}^{(j,\text{Amp})}\sigma^{(j)}_{x}$ (1)
×$\times\cos\left(\omega_{0}^{(j)}t+\frac{\Omega_{2}}{\Omega_{1}^{(j,\text{Phase})}}\sin\left(\Omega_{1}^{(j,\text{Phase})}t\right)\right)$
where $\omega_{0}^{(j)}$ is the $j$-th qubit’s transition frequency, $\sigma_{x}$ is a Pauli matrix, and $\Omega_{2}$ is a parameter quantifying the phase modulation amplitude. For this article we set $\hbar=1$ while all occurring frequencies are angular frequencies. $\Omega_{1}^{(j,\text{Amp})}$ is the Rabi frequency determined by the RF field amplitude for qubit $j$ and $\Omega_{1}^{(j,\text{Phase})}$ is set as a parameter to create the second dressing field for the qubits via phase modulation. For our gate scheme, we choose $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}},$ (optimization over this condition may result in longer coherence time [24]). To ensure $\Omega_{1}^{\text{Amp}}=\Omega_{1}^{\text{Phase}}$, a calibration of RF-amplitudes is carried out by recording and fitting Rabi oscillations of each individual qubit. The effect on the entangling gate in case of detuned Rabi frequencies is investigated in appendix F.
First, in Sec. III, we focus on the effect of the double dressing field according to Eq. (1) on individual qubits. Then, we describe the experimental procedure and results for the entangling gate in Sec. IV.
@@ III Double dressing of single qubits
The MAGIC scheme requires a magnetic sensitive qubit transition, making the qubits therefore susceptible to decoherence due to magnetic field fluctuations. We demonstrate experimentally that the double dressing field acts as continuous dynamical decoupling and improves the coherence time significantly, by suppressing the effect of addressing frequency fluctuations as well as amplitude fluctuations of the RF driving field. The idea of using double dressed states to counter noise has been established in previous works on nitrogen-vacancy centers in diamond [25, 23, 24]. Here, we extend double dressing to individual trapped ions, and investigate the coupling of the qubit resonance to vibrational modes by the phase modulated driving field.
[…]
@@
The double dressed scheme is illustrated for one qubit in Fig. 1 and is described in appendix A in detail. In brief, the ionic qubit in a harmonic oscillator potential is modeled as
$H=\frac{\omega_{0}}{2}\sigma_{z}+\nu b^{\dagger}b+\frac{\eta\nu}{2}\sigma_{z}\left(b+b^{\dagger}\right).$ (2)
Here, we include the COM vibrational mode and denote $b$ as the ladder operator of the motional quanta in the effective harmonic trapping potential. $\nu$ is the motional mode frequency and $\eta=g_{I}\mu_{B}(\partial_{z}B)\nu^{-3/2}/\sqrt{2Nm_{\text{Yb}}}$ is the effective Lamb-Dicke parameter proportional to the magnetic field gradient $(\partial_{z}B)$ [7]. $g_{I}$ is the effective $g$-factor in the electronic ground state of ${}^{171}\text{Yb}^{+}$, $m_{\text{Yb}}$ is the mass of ${}^{171}\text{Yb}$, $\mu_{B}$ is Bohr’s magneton, and $N$ is the number of ions in a linear Coulomb crystal. When turning on the driving field (Eq. (1)), we obtain the Hamiltonian,
@@
in the rotating frame w.r.t. $\omega_{0}/2\sigma_{z}$. Here, $S_{x}=\sigma_{z},S_{y}=\sigma_{y},S_{z}=-\sigma_{x}$ are the dressed basis operators. Dressing suppresses fluctuations of the bare state energy gap $\delta\omega_{0}\ll\Omega_{1}$ caused by either frequency fluctuations of the driving field, or by fluctuations of magnetically sensitive bare qubit states. Phase modulation of the dressing field, in addition, suppresses fluctuations in the amplitude of $\Omega_{1}$: Phase modulation creates, in an appropriate rotating frame, an effective on-resonance second drive in the dressed basis. This second drive dresses the qubit a second time, and thus dynamically decouples amplitude noise fluctuations in the drive $\delta\Omega_{1}$, which would otherwise reduce the coherence time of the dressed qubit. These fluctuations are suppressed as long as $\delta\Omega_{1}\ll\Omega_{2}$. Furthermore, undesired coupling of the qubit to the motional state, which leads to a decay in the Rabi oscillations, is suppressed as long as $\Omega_{2}\gg\eta\nu$ and $|\Omega_{1}-\nu|\ll\Omega_{2}\ll|\Omega_{1}+\nu|$ (see appendix A for a detailed description).
@@ III.1 Extending the coherence time
Figure 2: Relaxation time $T_{2,\rho}$ of a single qubit as a function of the modulation depth $\Omega_{2}/\Omega_{1}$. The qubit is initialized in state $|0\rangle$, which is a superposition state in the dressed state basis. The phase-modulated driving field is continuously applied and the coherence time $T_{2,\rho}$ is extracted from a exponential decay of the Rabi fringe contrast. At a modulation depth $\Omega_{2}/\Omega_{1}\approx 0.75$, the relaxation time is enhanced by two orders of magnitude compared to the scenario without second dressing field.
@@
The impact of the phase modulated dressing field on a qubit’s coherence time is shown in Fig. 2. We study the relaxation of a single qubit initialized in state $|0\rangle$, while the phase-modulated driving field given in Eq. (1) is continuously applied. Note, that this is an equal super-position state in the dressed basis, therefore, by fitting the decay of the oscillations’ contrast to an exponential decay we can extract the coherence time $T_{2,\rho}$ of the dressed state. For $\Omega_{2}/\Omega_{1}=0.75$, used for the entangling gate mechanism, we observe two orders of magnitude improvement in $T_{2,\rho}$, up to 0.27(2) s. The reduced coherence time for $\Omega_{2}/\Omega_{1}\approx 1$ is either explained by second-order fluctuations of $\delta\Omega_{2}$ [23] or by the quantum stark shift induced by the counter rotating terms.
III.2 Coupling to the motional degrees-of-freedom
The interaction between the qubits’ internal state is mediated via the ions’ common vibrational motion. Therefore, we investigate the qubit-phonon coupling, first with a single ion. We do this by probing a qubit’s coherence in a Ramsey-type experiment. An initial resonant $\pi/2$-pulse prepares the qubit in the eigenstate $\ket{+}$ of the driving field. Then the phase-modulated driving field is applied creating an initial spin-lock condition. A final $\pi/2$-pulse with variable phase projects the state vector into the $z-$basis, creating Ramsey fringes. Since the phase-modulated driving field is continuously applied, the state vector experiences continuous Rabi oscillations, traversing the Bloch sphere. Therefore, to map the coherence, we aim to apply the second Ramsey pulse while the state vector is in a coherent superposition on the Bloch sphere equator. We confirm the required evolution time by a projective measurement along the $z-$axis to ensure the Bloch vector being located in the equatorial plane of the Bloch sphere prior to Ramsey readout. Fig. 3 depicts the Ramsey fringe contrast measured as outlined above, for multiple steps in time up to the approximate entangling gate time of $300\leavevmode\nobreak\ \mathrm{\mu}s$, at $\Omega_{1}=2\pi\times$ 94.8 kHz and $\Omega_{1}=2\pi\times$ 61 kHz, respectively. In both cases, $\Omega_{2}/\Omega_{1}=\leavevmode\nobreak$0.75. At $\Omega_{1}=2\pi\times$ 94.8 kHz, close to the axial trap frequency, significant qubit-phonon entanglement at about 150 $\mu$s occurs, indicated by the reduction of the qubit’s internal coherence.

### v2

@@ Abstract
A novel two-qubit entangling gate for trapped-ion quantum processors is proposed theoretically and demonstrated experimentally. During the gate, double-dressed quantum states are created by applying a phase-modulated continuous driving field. The speed of this gate is an order of magnitude higher than that of previously demonstrated radio frequency (RF) controlled two-qubit entangling gates in static magnetic field gradients. At the same time, the field driving the gate, dynamically decouples the qubits from amplitude and frequency noise, increasing the qubits’ coherence time by three orders of magnitude. The gate requires only a single continuous RF field per qubit, making it well suited for scaling a quantum processor to large numbers of qubits. Implementing this entangling gate, we generate the Bell states |$\ket{\Phi^{+}}$ and |$\ket{\Psi^{+}}$ in $\leq 313$ $\mathrm{\mathrm{\mu}}$s with fidelities up to $98^{+2}_{-3}\leavevmode\nobreak$in a static magnetic gradient of only 19.09 T/m. At higher magnetic field gradients, the entangling gate speed can be further improved to match that of laser-based counterparts.
@@ I Introduction
Trapped atomic ions are among physical platforms well suited for quantum information processing [1, 2]. Outstanding performance has been realized in trapped-ion quantum processors by using focused lasers to coherently control individual ionic qubits [3, 4, 5, 6, 7, 8]. Trapped ions coherently controlled by radio frequency (RF) signals are particularly suited for scalability, since technological challenges associated with using laser light for coherent control of qubits are avoided in this laser-free approach [9, 10, 11, 12, 13]. High fidelity single- and two-qubit gates [14, 15, 16, 17, 18], as well as complete quantum algorithms [19, 20] have been realized with RF-controlled ions. Single-qubit rotations are performed with low cross-talk [21], which is crucial for fault-tolerant quantum computing.
Conditional quantum dynamics with trapped ion qubits using RF radiation requires a magnetic field gradient to couple internal qubit states to the ions’ motional states [9, 10, 22]. RF-based coherent control can be achieved either by far-field RF radiation applied in a local static magnetic field gradient [13, 15, 19], or by a dynamic magnetic field gradient [12, 4, 17]. While the two approaches share important features, for instance, laser-free coherent manipulation and magnetic field gradient induced coupling (MAGIC), each approach offers distinct advantages when it comes to scaling up ion-based quantum computing.
In both RF-based approaches, two-qubit gate speed has lagged behind its laser-based counterparts [17, 15]. Here, we introduce and experimentally realize a novel two-qubit entangling gate with gate speed an order of magnitude faster than previous RF-based gates in static magnetic field gradients [15] and a factor of two faster than previous two-qubit gates using the dynamic gradient approach [17]. Furthermore, in contrast to previous work, only a single phase modulated RF field per ion is required for its operation, further simplifying the scaling up of trapped-ion quantum computers.
RF-controlled qubits in static magnetic gradients often have a relatively short coherence time due to magnetic field fluctuations. Common practice to counteract decoherence is dynamical decoupling [23, 24]. Inspired by nuclear magnetic resonance refocusing pulses [25, 26, 27], refocusing techniques, referred to as pulsed dynamical decoupling, were designed for quantum registers to increase coherence time [28, 29, 30, 31, 32, 33, 34, 35]. The ideas of dynamical decoupling have also been extended to protect various gate operations by composite pulses [36, 37, 38, 39, 40]. Alternatively, dynamical decoupling is achieved by applying a continuous driving field, rather than pulses. Continuous dynamical decoupling, too, has been successfully incorporated into entangling gates in trapped ions [41, 42, 43, 15]. The gate presented in this work utilizes continuous dynamical decoupling to create and entangle double-dressed qubit states [44, 45], which offer built-in robustness against external noise making additional dynamical decoupling superfluous. Using these double-dressed qubits, we improved the coherence time by three orders-of-magnitude compared to the bare qubits.
The gate implemented here can be interpreted as a geometric phase gate [46, 47, 48, 49, 50, 51] between double-dressed qubits. A variant of double-dressed two-qubit gates has been theoretically proposed for spin qubits associated with silicon vacancy centers in diamond [52]. To our knowledge, the work presented here is the first experimental demonstration of a two-qubit gate based on double-dressed states in any physical platform.
This article is structured as follows: After a brief description of the experimental apparatus in Sec. II, we give an account of the experimental investigation of double dressing individually addressed qubits in a two-ion crystal in Sec. III. Then, the theoretical description of the entangling gate driven by a phase modulated field is presented in Sec. IV.1, before we describe the experimental procedure and results for this two-qubit entangling gate in Sec. IV.2 and IV.3, respectively. We conclude this article with a discussion of phase modulated vs. amplitude modulated driving fields for entangling gates as well as a summary of results and an outlook on future work in Sec. V.
@@ II Experimental overview
We investigate double dressed state dynamics of a single qubit and implement conditional quantum dynamics with two qubits using 171Yb+ ions trapped in a macroscopic linear Paul trap with radial and axial center-of-mass (COM) angular frequencies of $2\pi\times 380$ kHz and $\nu=2\pi\times 98.08$ kHz, respectively. The experimental setup is similar to [18] and only briefly described here. The two qubits are encoded in hyperfine states of the electronic ground state of 171Yb+ ions, $|0\rangle\equiv{|^{2}\rm S_{1/2},\rm F=0,\rm m_{\rm F}=0\rangle}$ and $|1\rangle\equiv{|^{2}\rm S_{1/2},\rm F=1,\rm m_{\rm F}=-1\rangle}$. A static magnetic field gradient of $19.09(1)$ T/m applied along the trap axis (z-axis) yields individual Zeeman-shifts of the two qubits and results in individual angular resonance frequencies, $\omega_{0}^{(1)}$ and $\omega_{0}^{(2)}$, of the magnetic dipole transition between qubit states. Both frequencies are near $2\pi\times 12.6$ GHz and differ by $2\pi\times 4.342(1)$ MHz. Thus, qubits are individually coherently controlled with low cross-talk using global RF radiation [21].
The entangling gate proposed and implemented here takes advantage of a state selective force induced by the static magnetic gradient field coupling the internal qubit states to the axial vibrational states of the 2-ion crystal [9]. The coupling strength between qubit states and motional states, the latter acting as a quantum bus between qubits [1], is given by the effective Lamb-Dicke parameter, $\eta=g_{F}\mu_{B}(\partial_{z}B)\nu^{-3/2}/\sqrt{2Nm_{\text{Yb}}}$ that is proportional to the magnetic field gradient, $\partial_{z}B$ [9], $g_{F}$ is the hyperfine Landé g-factor of ${}^{171}\text{Yb}^{+}$ (appendix A), $m_{\text{Yb}}$ is the mass of ${}^{171}\text{Yb}$, $\mu_{B}$ is Bohr’s magneton, and $N$ is the number of ions in a linear Coulomb crystal. In the experiments presented here $\eta=0.0329$. Therefore, when using magnetic gradient induced coupling (MAGIC), laser light is not required for implementing conditional gates with trapped ions.
The ions are cooled close to their motional ground state in two stages. Initial Doppler cooling is followed by RF sideband cooling of both vibrational modes giving a mean phonon number of 0.6(5) in the center-of-mass (COM) mode measured with sideband thermometry [53]. The heating rate of one trapped ion in the current experimental setup is 0.19(3) phonons/ms for this mode.
We use an Arbitrary Waveform Generator (AWG) to generate the phase-modulated RF driving fields that double dress both qubits [45, 54, 44].
@@ III Double dressing of single qubits
The MAGIC scheme in a static gradient field requires a magnetic sensitive qubit transition, making the qubits susceptible to decoherence due to magnetic field fluctuations. We demonstrate experimentally that the double dressing field acts as continuous dynamical decoupling and improves the coherence time significantly by suppressing the effect of addressing frequency fluctuations as well as amplitude fluctuations of the RF driving field. The idea of using double dressed states to counter noise has been established in previous works on single qubits in nitrogen-vacancy centers in diamond [55, 44, 56]. Here, we extend double dressing to individual trapped ions in a two-ion Coulomb crystal, and investigate the coupling of the qubit resonance to vibrational modes by the phase modulated driving field.
@@
The double dressed scheme is illustrated for one qubit in Fig. 1 and is described in appendix A in detail. In brief, the ionic qubit in a harmonic oscillator potential is modeled as [9]
$H=\frac{\omega_{0}^{(j)}}{2}\sigma_{z}+\nu b^{\dagger}b+\frac{\eta\nu}{2}\sigma_{z}\left(b+b^{\dagger}\right).$ (1)
Here, $\sigma_{z}$ is the the Pauli operator describing the qubit’s eigenstates. We consider the center-of-mass (COM) vibrational mode of two ions and denote $b$ as the ladder operator of the motional quanta in the effective harmonic trapping potential. In the mathematical description in this article we set $\hbar=1$. In all experiments reported in this article, we trap two ions (i.e., $N=2$) while the number of ions simultaneously addressed by a dedicated RF driving field changes. For the single-qubit experiments discussed in this section as well as for the two-qubit experiments in section IV, undesired crosstalk between qubits is negligible [21].
We drive one ion near-resonance with a phase modulated field such that
where $\Omega_{1}$ is the Rabi frequency of the RF drive, $\sigma_{x}$ a Pauli operator, and $\Omega_{2}$ is a parameter quantifying the phase modulation amplitude.
When the driving field is applied (Eq. (2)), we obtain the Hamiltonian,
@@
in the rotating frame w.r.t. $\omega_{0}\sigma_{z}/2+\Omega_{2}\sigma_{z}/2\cos\left(\Omega_{1}t\right)$. Here, $S_{x}=-\sigma_{z},S_{y}=\sigma_{y},S_{z}=\sigma_{x}$ are the dressed basis operators. Dressing suppresses fluctuations of the bare state energy gap $\delta\omega_{0}\ll\Omega_{1}$ caused by either frequency fluctuations of the driving field, or by fluctuations of magnetically sensitive bare qubit states which in turn may be caused by electric noise fields.
Phase modulation of the dressing field, in addition, suppresses fluctuations in the amplitude of $\Omega_{1}$: Phase modulation creates, in an appropriate rotating frame, an effective on-resonance second drive in the dressed basis. This second drive dresses the qubit a second time, and thus dynamically decouples amplitude noise fluctuations in the drive $\delta\Omega_{1}$, which would otherwise reduce the coherence time of the dressed qubit (Fig. 1). These fluctuations are suppressed as long as $\delta\Omega_{1}\ll\Omega_{2}$. Furthermore, undesired coupling of the qubit to the motional state, which leads to a decay in the Rabi oscillations, is suppressed as long as $\Omega_{2}\gg\eta\nu$ and $|\Omega_{1}-\nu|\ll\Omega_{2}\ll|\Omega_{1}+\nu|$ (appendix A gives a detailed description).
@@ III.1 Extending the coherence time
Figure 2: Relaxation time $T_{2,\rho}$ of a single qubit as a function of the modulation depth $\Omega_{2}/\Omega_{1}$. The qubit is initialized in state $|0\rangle$, which is a superposition state in the dressed state basis. The phase-modulated driving field is continuously applied and the coherence time $T_{2,\rho}$ is extracted from an exponential decay of the Rabi fringe contrast. At a modulation depth $\Omega_{2}/\Omega_{1}\approx 0.75$, the relaxation time is enhanced by three orders of magnitude compared to the scenario without second dressing field.
@@
The impact of the phase modulated dressing field on a qubit’s coherence time is shown in Fig. 2. We study the relaxation of a single qubit in a two-qubit crystal initialized in state $|0\rangle$, while the phase-modulated driving field given in Eq. (2) is continuously applied. Note, that this is an equal super-position state in the dressed basis, therefore, by fitting the decay of the oscillations’ contrast to an exponential decay we can extract the coherence time, $T_{2,\rho}$ of the dressed state.
When no modulation is applied, we observe ordinary Rabi oscillations (not shown) with frequency $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$kHz, close to the trap frequency $\nu=2\pi\times 98.08\leavevmode\nobreak$kHz. At low modulation strength, the coherence time is mainly limited by the near-resonant coupling of the qubit to the COM motional mode. With increasing modulation amplitude, $\Omega_{2}/\Omega_{1}$, the coherence time of the dressed state, $T_{2,\rho}$ increases.
The diminished coherence time at $\Omega_{2}/\Omega_{1}\approx 0.1$ is attributed to the quantum Stark shift. For $\Omega_{1}=2\pi\times 94.8\leavevmode\nobreak$ kHz, $\Omega_{2}/2$ is nearly equal to $\nu\eta\leavevmode\nobreak$ kHz, and motional states are shifted close to resonance with phase modulation that drives the transition between dressed states (details are given in appendix B).
For $\Omega_{2}/\Omega_{1}=0.75$, used for the entangling gate mechanism in this work, we observe three orders of magnitude improvement in $T_{2,\rho}$, up to 0.27(2) s. The slightly reduced coherence time for $\Omega_{2}/\Omega_{1}\approx 1$ is due to second-order fluctuations of $\delta\Omega_{2}$ [44] and to the quantum stark shift induced by counter rotating terms.
III.2 Coherent coupling to motional degrees-of-freedom
The interaction between the qubits’ internal state is mediated via the ions’ common vibrational motion. Therefore, we investigate the qubit-phonon coupling in more detail by probing a qubit’s coherence in a Ramsey-type experiment when addressing a single ion with the phase-modulated driving field.
First, a resonant and not phase-modulated $\pi/2$-pulse prepares the qubit in the eigenstate |$\ket{+}=\frac{1}{\sqrt{2}}\left(\ket{0}+\ket{1}\right)$ of the driving field. Then, the phase-modulated driving field is applied. A final $\pi/2$-pulse with variable phase projects the state vector into the $z-$basis, creating Ramsey fringes.
Since the phase-modulated driving field is continuously applied, the state vector experiences continuous Rabi oscillations, traversing the Bloch sphere. Therefore, to map the coherence, we apply the second Ramsey pulse when the state vector is in a coherent superposition on the Bloch sphere equator. We confirm the required evolution time by a projective measurement along the $z-$axis to ensure the Bloch vector is located in the equatorial plane of the Bloch sphere prior to Ramsey readout. By this, we ensure that we choose interaction times such that the dressed basis coincides with the bare basis, that is, when the rotating frame of Eq. (3) concurs with the lab frame.
Fig. 3 depicts the Ramsey fringe contrast measured as outlined above, for multiple steps in time at $\Omega_{1}=2\pi\times$ 94.8 kHz (blue data points) and $\Omega_{1}=2\pi\times$ 61 kHz (red data points), respectively. In both cases, $\Omega_{2}/\Omega_{1}=\leavevmode\nobreak$0.75. At $\Omega_{1}=2\pi\times$ 94.8 kHz, close to the axial trap frequency, significant qubit-phonon entanglement at about 150 $\mu$s occurs, indicated by the reduction of the qubit’s internal coherence.